OpenStudy (alexa):

A perpendicular bisector to AB is drawn through point C and extended to point D. Arrange the pairs of points in order of the y-intercepts of their corresponding perpendicular bisectors, starting with the smallest and ending with the largest. A(-4,5) and B (8,9) A(2,4) and B (-8,6) A(5,4) and B (7,2) A(2,9) and B (-4,3) A(3,-2) and B (9,-12) A(4,10) and B (8,12)

OpenStudy (prathamesh_m):

To find perpendicular bisector first find the midpoint of the points A and B by using the section formula. Then find the slope of the perpendicular bisector(You can find the slope of line segment AB). Then using the slope of the perpendicular bisector and the midpoint of AB(which lies on it) u can find its equation. Then it should be pretty straightforward to find their y-intercepts.

OpenStudy (supersmart1001):

so what's the slope for line AB?

OpenStudy (alexa):

ok ok im here

OpenStudy (supersmart1001):

ok so what's the slope for line AB?

OpenStudy (supersmart1001):

I got 1/3 when I solved it.

OpenStudy (supersmart1001):

did u get that

OpenStudy (alexa):

how did you get 1/3 because when you subtract 8 minus -4 isn't it 4 or somethin?

OpenStudy (prathamesh_m):

its 12 8-(-4)=12

OpenStudy (prathamesh_m):

9-5=4 so slope is 1/3

OpenStudy (alexa):

oh its ok. OH so after i find all the slopes what do i do after that?

OpenStudy (supersmart1001):

good, so now, you know the slope for your line now we just need to plug that into point-slope form to get the eq for the line

OpenStudy (supersmart1001):

now we are at Y=1/3x+b right? Now we are solving for the Y-intercept?

OpenStudy (alexa):

yea, the y-intercept

OpenStudy (supersmart1001):

so now just solve for y and you'll get your intercept.

OpenStudy (alexa):

how would i do that?

OpenStudy (supersmart1001):

you have to distribute the 1/3 first

OpenStudy (supersmart1001):

distribute the 1/3 to the 4

OpenStudy (supersmart1001):

how do you get the 5 to the other side?

OpenStudy (alexa):

subtraction?

OpenStudy (supersmart1001):

I got to Y-5=1/3x+4/3

OpenStudy (alexa):

how did you get the 4/3? I just want to know how to work out the first one so i get the others correct.

OpenStudy (supersmart1001):

srry i need some time to solve this im getting confused

OpenStudy (supersmart1001):

@phi

OpenStudy (supersmart1001):

@Sachintha

OpenStudy (alexa):

its ok don't worry.

OpenStudy (supersmart1001):

ok let me see if anyone else can help

OpenStudy (supersmart1001):

@Vocaloid

OpenStudy (supersmart1001):

@tigerlover

OpenStudy (supersmart1001):

@nickj123

OpenStudy (supersmart1001):

@jtug6

OpenStudy (alexa):

alright but thanks for helping! :)

OpenStudy (supersmart1001):

hope these guys and ladys can help srry again that i cant help m8

OpenStudy (alexa):

This is for Plato, you need to miss only one

OpenStudy (phi):

we want the equation of the *perpendicular bisector* i.e. perpendicular to the line segment AB if the slope of AB is m the slope of the perpendicular is -1/m ("flip" the slope, then multiply by -1) for example, for the first A,B pair A(-4,5) and B (8,9) you get a slope m = 1/3 now "flip" that fraction to get 3/1 or 3 then multiply by -1 to get -3 -3 is the slope of the perpendicular.

OpenStudy (phi):

next, you need a point on the perpendicular bisector. by definition, one point is the point exactly in between A and B to find that point, I would find the average of the x value of A and the x value of B and do the same for the y values.

OpenStudy (phi):

For example, with A(-4,5) and B (8,9) the x value of the midpoint is (-4+8)/2 = 4/2 = 2 the y value is (5+9)/2 = 14/2 = 7 so (2,7) is a point on the perpendicular bisector. we can now find the equation of the line of this perpendicular bisector. one way is write y = m x + b fill in m (which is -3) y = -3 x + b now fill in x and y for a point on the line. in other words, we use (2,7): 7 = -3*2 + b 7 = -6+b add 6 to both sides: 6+7 = 6-6+b 13= b that means the y-intercept is 13 finally, we have the equation. y = -3x + 13 so the first pair has a y-intercept of 13 we have to do all that work for each pair of A, B. (Gack!)

OpenStudy (sachintha):

@phi I am trying to form a general equation for the perpendicular bisector. But, I am getting the wrong equation. -_-

OpenStudy (phi):

I got a messy one. It's hardly worth using \[b= \frac{ (A_x^2 +A_y^2) - (B_x^2 +B_y^2)}{2(A_y-B_y)} \]

OpenStudy (alexa):

which formula do i use? Its match and match so if i get one wrong it messes up everything

OpenStudy (phi):

I would first: find the slope of each A/B pair. can you do that ?

OpenStudy (alexa):

ok ill try,gonna take me some time

OpenStudy (alexa):

1. 1/3 2 -10/2 3 2/-2 4. -6/-6 6.4/2 5. 6/-10

OpenStudy (alexa):

is that right? Can someone confirm

OpenStudy (phi):

almost. it looks like you did change in x divided by change in y which is upside down.

OpenStudy (phi):

this problem has to be done carefully. so please fix those slopes.

OpenStudy (sachintha):

 Point P is an arbitrary position on the perpendicular bisector and C is the mid point of A and B. Gradient of line AB = \(\Large\frac{y_2-y_1}{x_2-x_1}\) Gradient of line PC = \(-\Large\frac{x_2-x_1}{y_2-y_1}\) I got the equation of line PC as, \(y-y_3=-\Large\frac{x_2-x_1}{y_2-y_1}\normalsize(x_3-x)\) I can't figure out what I have done wrong. >_<